OpenStudy (anonymous):

HELP!!!! Center at the origin, Transverse axis at x-axis, distance between foci is 2√97, latus rectum is 32/9

OpenStudy (anonymous):

HELP PLEASE!!!

OpenStudy (tkhunny):

So, it's an ellipse. Center at the Origin gives: \(\dfrac{x^2}{a^{2}} + \dfrac{y^{2}}{b^{2}} = 1\) x-Axis is Transverse (or Major) Axis give: a > b Distance between foci gives \(c = \sqrt{97}\) Latus Rectum is 32/9 gives \(2\cdot\dfrac{b^{2}}{a} = 32/9\) Well, that's all the problem statement gave us. Now what?

OpenStudy (anonymous):

its hyperbola.. hehe wait, 2√97, what happened to 2?

OpenStudy (tkhunny):

Well, you shoudl have said so! c is HALF the distance between the foci. Just like a is half the length fo the Transverse Axis and b is half the length of the congugate axis.

OpenStudy (anonymous):

sorry.. thank you :)

OpenStudy (tkhunny):

Change it to -y^2 instead of +y^2, and we're still on the right track.

OpenStudy (anonymous):

why is it changed to negative?

OpenStudy (tkhunny):

Because it an hyperbola and not an ellipse.

OpenStudy (anonymous):

yes yes.

OpenStudy (anonymous):

ahm, how to get a and b?